# Hierarchical Text Classification  (HTC) vs. eXtreme Multilabel Classification (XML): Two Sides of the Same Medal

### v2

@@ 5.2. HTC and XML methods on XML datasets
Due to resource limitations, HBGL fails on the Amazon 670k dataset and no results could be produced for HGCLR, too. The latter is because HGCLR faces several critical challenges that need to be addressed. The main problem is that the model uses numerous $O(n^{2})$ complexity functions in different parts of the code, namely in graph encoding and contrastive learning. Here, the inclusion of the extreme number of labels used in XML datasets is likely to cause memory problems on the GPU. With the help of the authors of HGCLR, the problem was narrowed down to HGCLR’s dense attention functions, which are difficult to optimize for large numbers of labels. For this reason, we were unable to complete experiments with HGCLR on XML datasets.
6. Discussion
Section 6.1 discusses the main findings, followed by Section 6.2, which addresses potential threats to the validity of the results. Section 6.3 examines how the results can be generalized in a broader NLP context.
6.1. Key Scientific Insights
Top Performers
Our main takeaway is that while XML methods are well suited for transfer to the HTC domain and achieve competitive results on HTC datasets, HTC methods are not able to manage the sheer size of labels in the XML datasets and are not suitable for use in the XML domain. XML models and their label-clustering techniques work well on HTC datasets and, due to their inherent method of creating an artificial hierarchy, do not even rely on the taxonomies that come with HTC datasets. On the HTC datasets NYT and RCV1-V2, the result of CascadeXML and HGCLR are very close together in first and second place for the P@1 to P@5 metrics and R-Precision, with CascadeXML slightly better on NYT and HGCLR narrowly winning on RCV1-V2.
The only HTC dataset, on which both HTC methods HBGL and HGCLR remain best on the R-Precision metric is the WoS dataset. This is likely due to the unique anatomy of this dataset. While all the other datasets have variable numbers of relevant labels for each document, with an average number of relevant labels of 3.2 and upwards, the WoS dataset contains exactly two relevant labels for each data point and also comes with a specific label hierarchy. Every document is assigned a single parent label and a single leaf label. This makes it more similar to single-label datasets than to multi-label datasets, and it may not be suitable for the clustering techniques used in XML models. It should also be noted that HGCLR outperforms HBGL on all three datasets from the HTC domain. At first glance, this seems to contradict the results presented by the authors of HBGL, which indicate that HBGL is the stronger model of the two, but can be explained by the different selection of performance metrics in the HBGL paper and this one. HBGL produces better F1 scores and HGCLR is stronger on the P@k metric and R-Precision. This means that HGCLR is better at identifying the top labels, but struggles to correctly identify all the relevant labels, resulting in worse F1 scores. HBGL seems to be the opposite. While its top predictions are not as strong as those of HGCLR or CascadeXML, it is better at correctly identifying relevant labels that are less obvious.
For the datasets selected from the XML domain, the comparison is not close. Neither HBGL nor HGCLR produce competitive results. HGCLR is unable to complete the training process on any of the XML datasets due to its $O(n^{2})$ complexity, where $n$ is the number of possible labels. HBGL faces similar problems and does not produce results on the Amazon-670K dataset, even on the $30k$ subset of training samples. For comparison, the largest HTC dataset regarding label space is NYT with $166$ labels. Even the smallest XML dataset in terms of labels, AmazonCat-13K, already contains $13,330$ labels. While HBGL is able to complete its training and testing pipelines on the smallest datasets from the XML world - AmazonCat-13K and Wiki10-31K - it still performs considerably worse than the models from the XML domain. Between these two, XR-Transformer wins the comparison on the Wiki10-31K and Amazon-670K datasets, and CascadeXML performs best on AmazonCat-13K.
Clustering Techniques
While both the HTC and XML methods belong to the hierarchical multi-label text classification domain, which leverages relational information about the label-space, their respective approaches to utilizing this information are quite different. The hierarchies employed by HTC datasets and methods differ from the HLTs created prior to the training process of XML models. While the former are based on semantic similarities, the latter are characterized by clustering labels together that have similar relevance frequencies and are relevant on documents with similar TF-IDF feature representations. The results of our experiments indicate that the former approach, which is oriented towards a human understanding of label similarity, may not be as beneficial for a machine-driven learning process as the latter. Furthermore, XML models are particularly strong on the P@1 metric, which could also be due to this difference in clustering strategy. Frequency-based clustering encourages the models to predict frequent labels even more often, which is beneficial for predicting head labels and increases P@1, especially when the evaluated dataset contains strong head labels that occur disproportionately, as is the case in the RCV1-V2 and NYT datasets. This is also supported by the results of HBGL** where the native taxonomy was replaced by an XML-style one. Although the improvements made are marginal and too small to claim that this version of clustering is superior in a broader machine-learning context, they are best in P@1.
Metrics
As mentioned in Section 4.5, the XML and HTC domains use different measures to evaluate model performance. In the domain of HTC, F1 scores are mainly used, and in the domain of XML, the P@k metric is predominant. However, the results of our experiments suggest that it is not sufficient to use only one or the other to capture all aspects of model performance. Models should ideally be evaluated on all of these metrics, and possibly even the R-Precision metric.
As presented by the authors of HBGL, HBGL outperforms HGCLR on both the Micro-F1 and Macro-F1 metrics, but in our experiments, HGCLR achieved better results on the P@k and R-Precision metrics, suggesting that HBGL is not stronger per se, but that the two models are rather good at different aspects of text classification. A comparison based on just one set of these metrics would have missed this observation. It would therefore be beneficial to make a comparison using a wider range of metrics, rather than focusing on just one set of them.
It should be noted that some models would need to be adapted to support both families of performance measures, as F1-scores are based on yes or no classifications and both P@k and R-Precision scores are based on rankings. The system under test would have to be able to support both paradigms.
6.2. Threat to Validity
There are a number of threats to validity that need to be discussed. The first relates to the structure of some of the datasets used. In some datasets, texts are labeled with arbitrary labels, that have no significant semantic meaning (i. e., ”!”, ”!!” , ”AE04752”). While the lack of semantic context in these labels may raise concerns about the overall suitability of such datasets for hierarchical classification, it is unlikely a problem for our purposes. Some of the evaluated models, like HBGL, can perform better by incorporating label semantics (using pre-trained embeddings of the textual representation of the label). Ablation studies for HBGL (Jiang et al., 2022) and HGCLR (Wang et al., 2022a) show that the impact of label semantics is relatively small.
Like the model examinations mentioned in Section 2, the datasets used in this study are exclusively in the English language. Consequently, the results presented in this work are solely applicable to the English language domain and may not be directly transferable to other languages.
Detailed hyperparameter optimization, e. g., in the form of a grid search, was out of scope due to the large number of experiments. Instead, we used the hyperparameters tuned by the authors of the respective methods. For the transfer from HTC to XML, we used the hyperparameters of the datasets that came closest to those of the other domain. These were mostly those from Wiki10-31k in the case of XML datasets for HTC methods and those from RCV1-V2 (HBGL configuration) in the case of HTC datasets for XML methods.
Furthermore, the dataset transfer between XML and HTC methods may not be optimal due to differences in their hierarchical structures. XML datasets do not inherently contain hierarchical labels, instead, they generate them through clustering. These clusters include meta labels, which guide the learning process towards the actual labels used for classification - the leaf nodes. HTC methods then apply these generated hierarchies in their models. However, HTC methods expect the entire hierarchy to consist of labels valid for classification. Since only the leaf nodes are actual labels and the rest are meta labels, HTC models are limited to classifying only at the final level of the hierarchy. This restriction may result in diminished effectiveness, as the model cannot classify at each hierarchical level. In other words, the XML datasets are actually only multi-label classification tasks, where the generated hierarchical label tree is a tool and not used to classify the documents on different levels.
Some datasets commonly used in the XML world, such as Wiki-500K and Amazon-3M, were included in the experimental setting in the first place but could not be tested for XR-Transformer and HTC methods due to resource limitations. We ran our experiments on NVIDIA H100 80GB GPUs with 1TB of RAM available but encountered OOM errors when training XR-Transformer on Wiki-500K and Amazon-3M. For the HTC methods, training time was the limiting factor. We were unable to complete the training within the 48-hour window imposed by our computing resource provider.
The performance results reported for CascadeXML by the authors of the model, Kharbanda et al. (2022), on the datasets Wiki10-31k and Amazon-670k could not be achieved. On the Wiki10-31k dataset, our models exhibit a mean deficit of $3.18$ points across the P@1, P@3, and P@5 metrics. On the Amazon-670k dataset, the analogous mean deficit is $3.76$ points. This might be due to the fact, that in contrast to our experiments, CascadeXML was run in an ensemble setting by the authors of the original paper, which can increase performance. On AmazonCat-13K, however, our numbers of CascadeXML are actually slightly better than the ensemble setting. The influence of the choice of language model(s) used and the impact of using ensemble should be analyzed in future work. The principal outcome of our analysis comparing HTC versus XML methods on the respective datasets is not affected because the CascadeXML used here without ensemble outperforms XR-Transformer on all HTC datasets and the AmazonCat-13k dataset.
6.3. Generalization
Our evaluation results on HTC datasets show that XML methods, in particular CascadeXML, are very competitive and can be considered for hierarchical text classification tasks. CascadeXML performed particularly well on the NYT dataset, even outperforming state-of-the-art HTC methods such as HGCLR and HBGL on some datasets. This suggests that the structure of the NYT dataset, which involves complex and deep hierarchical categorizations of news articles, is particularly well suited to the strengths of XML methods. At the same time, despite the challenging comparison, HGCLR has shown that HTC methods remain relevant and effective in certain scenarios. The results suggest that HTC methods could benefit from adaptation to larger datasets, which would require model refinements to cope with increasing complexity.
7. Conclusion
We investigated the cross-domain capabilities of Hierarchical Text Classification (HTC) and eXtreme Multi-Label Classification (XML) models, using cross-dataset transfer to evaluate their performance on unfamiliar datasets and label spaces. Our results show that XML models, in particular CascadeXML, demonstrate considerable adaptability and competitive accuracy when applied to HTC datasets, effectively handling smaller hierarchical structures despite being designed for large, unstructured label pools. However, the HTC models struggled with the sizes and sparsity of the XML datasets, suggesting scaling limitations.
Our results call for a rethinking of common clustering techniques for XML and HTC methods. Label clustering based on the features of the documents to which they are relevant could potentially be replaced or combined with a label frequency-based clustering approach. Other clustering techniques, such as neural clustering, could also be explored. Lastly, building on the initial success of XML models in the HTC world, extending the application of XML methods to other domains could be beneficial.
Acknowledgements.
Appendix
Appendix A Level-wise Number of Labels in the HTC Datasets
Table 5 shows the details of the hierarchical structure of the HTC datasets.
Table 5. Absolute number of labels in each hierarchy layer of the HTC datasets. Notice that only HTC datasets provide an inherent hierarchical label structure, while XML methods compute the label tree via clustering.
Dataset L1 L2 L3 L4 L5 L6 L7 L8
WoS 7 134
NYT 4 27 51 47 17 12 6 2
RCV1-v2 4 55 43 1
Appendix B Results for Leaf Runs
@@
The transfer of HTC datasets to the XML world can be achieved in a number of ways. In the HTC world, datasets come with semantic label hierarchies, and if a label at the leaf level of that hierarchy is relevant to a query, this means that all parent labels of that leaf label are implicitly relevant to that query as well. In our main experiments, we included both leaf labels and all implicit labels in the transfer of HTC datasets to the XML domain. In this additional analysis, we explore the other possible way of transferring the dataset, which ignores the implied positive labels derived from the hierarchy and focuses only on those labels that have no further children. To enable a fair comparison between XML and HTC models, the datasets were also adjusted for the HTC world, including removing all positive labels that are not leaf labels for each data point. It should be noted that this adjusted task is much harder for models from both worlds for two reasons. First, the number of relevant labels for each query is lower in this version, which has the consequence that P@k results are naturally worse. Second, information is lost in the discarding of labels that are not leaf labels in the hierarchy. The results of these runs with modified datasets can be seen in the following tables. On the NYT dataset, CascadeXML and HGCLR remain the best-performing models, with HBGL’s performance dropping strongly from $75.93$ to $42.30$ points in R-Precision compared to the main analysis. On the Web of Science dataset, which is a single-label dataset after the removal of all parent labels, HBGL overtakes HGCLR as the best-performing model with an R-Precision of $80.65$, narrowly beating out CascadeXML with an R-Precision of $77.47$. XR-Transformer achieves $69.97$ points and HGCLR is in last place with only $22.74$ points. On the RCV1-V2 dataset, HGCLR performs best again followed by CascadeXML and XR-Transformer. HBGL is not able to handle this version of the RCV1 dataset well and achieves an R-Precision of only $8.94$ points. Overall the results of HTC models on these adjusted datasets are very inconsistent. This is likely due to their reliance on relational information between labels. Part of this information is lost when only taking leaf labels into account in training and testing. XML models do not seem to rely on this information as heavily.
Table 6. New York Times Annotated Corpus
@@
Appendix C Additional Results for Wiki10-31K and Amazon-670K
Table 9 shows additional results for the Wiki10-31K dataset regarding P@10 and P@20, since the dataset has a higher average number of labels for sample. The general trend of the P@k metric continues also for higher values of $K$. Table 10 shows additionally the P@10 scores for Amazon-670K.
Table 9. Additional results for Wiki10-31K. For methods marked with † no results could be obtained due to resource limitations.
Wiki10-31K
Method P@1 P@2 P@3 P@5 P@10 P@20 R-Prec
HGCLR - - - - - - -
HBGL 53.10 52.79 53.15 52.32 42.95 27.71 26.83
XR-Transformer 88.12 84.87 80.39 70.74 53.82 36.76 38.93
CascadeXML 86.80 81.90 76.70 67.04 50.07 33.67 35.76
Table 10. Additional results for Amazon-670K. For methods marked with † no results could be obtained due to resource limitations.
Amazon-670K
Method P@1 P@2 P@3 P@5 P@10 R-Prec
HGCLR - - - - - -
HBGL - - - - - -
XR-Transformer 49.17 46.12 43.75 39.91 26.83 39.91
CascadeXML 47.86 45.10 42.88 39.15 26.19 39.66

### v3

@@ 5.2 HTC and XML methods on XML datasets
Due to resource limitations, HBGL fails on the Amazon 670k dataset and no results could be produced for HGCLR, too. The latter is because HGCLR faces several critical challenges that need to be addressed. The main problem is that the model uses numerous $O(n^{2})$ complexity functions in different parts of the code, namely in graph encoding and contrastive learning. Here, the inclusion of the extreme number of labels used in XML datasets is likely to cause memory problems on the GPU. With the help of the authors of HGCLR, the problem was narrowed down to HGCLR’s dense attention functions, which are difficult to optimize for large numbers of labels, which is why experiments with HGCLR on XML datasets did not complete.
5.3 Ablation: Results of the Leaf-Runs using XML models on HTC datasets
@@
The transfer of HTC datasets to the XML world can be achieved in a number of ways. In the HTC world, datasets come with semantic label hierarchies, and if a label at the leaf level of that hierarchy is relevant to a query, this means that all parent labels of that leaf label are implicitly relevant to that query as well. In our main experiments, we included both leaf labels and all implicit labels in the transfer of HTC datasets to the XML domain. In this additional analysis, we explore the other possible way of transferring the dataset, which ignores the implied positive labels derived from the hierarchy and focuses only on those labels that have no further children. To enable a fair comparison between XML and HTC models, the datasets were also adjusted for the HTC world, including removing all positive labels that are not leaf labels for each data point. It should be noted that this adjusted task is much harder for models from both worlds for two reasons. First, the number of relevant labels for each query is lower in this version, which has the consequence that P@k results are naturally worse. Second, information is lost in the discarding of labels that are not leaf labels in the hierarchy.
The results of these runs with modified datasets can be seen in Table 5. On the NYT dataset, CascadeXML and HGCLR remain the best-performing models, with HBGL’s performance dropping strongly from $75.93$ to $42.30$ points in R-Precision compared to the main analysis. On the Web of Science dataset, which is a single-label dataset after the removal of all parent labels, HBGL overtakes HGCLR as the best-performing model with an R-Precision of $80.65$, narrowly beating CascadeXML with an R-Precision of $77.47$. XR-Transformer achieves $69.97$ points, and HGCLR is in last place with only $22.74$ points. On the RCV1-V2 dataset, HGCLR performs best again, followed by CascadeXML and XR-Transformer. HBGL is not able to handle this version of the RCV1 dataset well and achieves an R-Precision of only $8.94$ points.
Table 5: Results of the Leaf-Runs using the XML models on the HTC datasets
@@
Overall, the results of HTC models on these adjusted datasets are very inconsistent. This is likely due to their reliance on relational information between labels. Part of this information is lost when only taking leaf labels into account in training and testing. XML models do not seem to rely on this information as heavily.
6 Discussion
Key Insights
Our main takeaway is that XML methods are well suited for transfer to the HTC domain and achieve competitive results on HTC datasets. However, HTC methods are not able to manage the sheer size of labels in the XML datasets (see Section 5.2). XML models and their label-clustering techniques work well on HTC datasets and, due to their inherent method of creating an artificial hierarchy, do not even rely on the taxonomies that come with HTC datasets. On the HTC datasets NYT and RCV1-V2, the result of CascadeXML and HGCLR are very close together in first and second place for the P@1 to P@5 metrics and R-Precision, with CascadeXML slightly better on NYT and HGCLR narrowly winning on RCV1-V2. In other words, a generic (extreme) multi-label classification method is better at hierarchical text classification than methods specifically designed for hierarchical text classification.
The only HTC dataset on which both HTC methods HBGL and HGCLR remain best on the R-Precision metric is the WoS dataset. This is likely due to the unique anatomy of this dataset. While all the other datasets have variable numbers of relevant labels for each document, see Table 2, the WoS dataset contains exactly two relevant labels for each data point and also comes with a specific label hierarchy: Every document is assigned a single parent label and a single leaf label. This makes it more similar to single-label datasets than to multi-label datasets, and it may not be suitable for the clustering techniques used in XML models.
It should also be noted that HGCLR outperforms HBGL on all three datasets from the HTC domain. At first glance, this seems to contradict the results presented by the authors of HBGL, which indicate that HBGL is the stronger model of the two, but this can be explained by the different selection of performance metrics in the HBGL paper and this one. HBGL produces better F1 scores, and HGCLR is stronger on the P@k metric and R-Precision. This means that HGCLR is better at identifying the top labels but struggles to correctly identify all the relevant labels, resulting in worse F1 scores. HBGL seems to be the opposite: Although its top predictions are not as strong as those of HGCLR or CascadeXML, it is better at correctly identifying relevant labels that are less obvious.
For the XML datasets selected from the XML domain, the comparison is not close. Neither HBGL nor HGCLR produce competitive results. HGCLR is unable to complete the training process on any of the XML datasets due to its $O(n^{2})$ complexity, where $n$ is the number of possible labels. HBGL faces similar problems and does not produce results on the Amazon-670K dataset, even on a $30k$ subset of training samples. For comparison, the largest HTC dataset regarding label space is NYT with $166$ labels. Even the smallest XML dataset in terms of labels, AmazonCat-13K, already contains $13,330$ labels. While HBGL is able to complete its training and testing pipelines on the smallest datasets from the XML world, AmazonCat-13K and Wiki10-31K, it still performs considerably worse than the models from the XML domain (see Section 5.2). Between these two, XR-Transformer is better on the Wiki10-31K and Amazon-670K datasets, and CascadeXML performs best on AmazonCat-13K.
Partitioning the Label Space
Both the HTC and XML methods aim to tackle multi-label text classification and make use of a hierarchy to leverage relational information about the label space. However, their approaches to utilizing this information are quite different. The hierarchies employed by HTC datasets and methods differ from the hierarchical label trees created prior to the training process of XML models. While the former is based on semantic similarities, the latter are characterized by clustering labels together that have similar relevance frequencies and are relevant on documents with similar TF-IDF feature representations. The results of our experiments indicate that the former approach, which is oriented towards a human understanding of label similarity, may not be as beneficial for a machine-driven learning process as the latter. Furthermore, XML models are particularly strong on the P@1 metric, which could also be due to this difference in clustering strategy. Frequency-based clustering encourages the models to predict frequent labels even more often, which is beneficial for predicting head labels and increases P@1, especially when the evaluated dataset contains strong head labels that occur disproportionately, as is the case in the RCV1-V2 and NYT datasets. This is also supported by the results of HBGL** where the native taxonomy was replaced by an XML-style one. Although the improvements are too marginal to claim that this version of clustering is generally superior, they are best in P@1.
Metrics
As mentioned in Section 3.2, the XML and HTC domains use different measures to evaluate model performance. In the domain of HTC, F1 scores are mainly used, and in the domain of XML, the P@k metric is predominant. However, the results of our experiments suggest that it is not sufficient to use only one or the other to capture all aspects of model performance. Models should ideally be evaluated on all of these metrics and possibly even the R-Precision metric.
As presented by the authors of HBGL, HBGL outperforms HGCLR on both the Micro-F1 and Macro-F1 metrics, but in our experiments, HGCLR achieved better results on the P@k and R-Precision metrics, suggesting that HBGL is not stronger than HGCLR per se, but that the two models are rather good at different aspects of text classification. A comparison based on just one set of these metrics would have missed this observation. It would, therefore, be beneficial to make comparisons using a wider range of metrics. Note that some models would need to be adapted to support both families of performance measures, as F1-scores are based on yes or no classifications, and both P@k and R-Precision scores are based on rankings. The system under test would have to be able to support both paradigms (see Section 3.2).
Threats to Validity
There are a number of threats to validity that need to be discussed. The first relates to the structure of some of the datasets used. The dataset transfer between XML and HTC methods may not be optimal due to differences in their hierarchical structures. XML datasets do not inherently contain hierarchical labels. Instead, they are generated through clustering. These clusters include meta labels, which guide the learning process towards the actual labels used for classification, the leaf nodes. If we apply HTC methods on such generated hierarchies, as in Appendices 3.3 and 5.2, the HTC methods expect the entire hierarchy to consist of labels valid for classification. Since only the leaf nodes are actual labels and the rest are meta labels, HTC models are limited to classifying only at the final level of the hierarchy. In other words, the XML datasets are actually only multi-label classification tasks, where the generated hierarchical label tree is a tool and not used to classify the documents.
The performance of CascadeXML reported by Kharbanda et al. [16] on Wiki10-31k and Amazon-670k could not be achieved. On the Wiki10-31k dataset, our models exhibit a mean deficit of $3.18$ points across the P@1, P@3, and P@5 metrics. On the Amazon-670k dataset, it is $3.76$ points. This might be due to CascadeXML being run in an ensemble setting while we employ the BERT-only variant. However, on AmazonCat-13K our numbers of CascadeXML are actually better than the ensemble. The outcome of our analysis comparing HTC versus XML methods on the respective datasets is not affected because CascadeXML with BERT outperforms XR-Transformer on all HTC datasets and the AmazonCat-13k dataset.
Detailed hyperparameter optimization, e. g., in the form of a grid search, was out of scope due to the number of experiments. Instead, we used the hyperparameters tuned by the authors of the respective methods. For the transfer from HTC to XML, we used the hyperparameters of the datasets that came closest to those of the other domain. These were mostly those from Wiki10-31k in the case of XML datasets for HTC methods and those from RCV1-V2 (HBGL configuration) in the case of HTC datasets for XML methods.
In summary, our results show that XML methods, in particular CascadeXML, are very competitive and can be considered for hierarchical text classification tasks. CascadeXML works particularly well on the NYT dataset, even outperforming state-of-the-art HTC methods such as HGCLR and HBGL. This suggests that the structure of the NYT dataset, which involves complex and deep hierarchical categorizations of news articles, is particularly well suited to the strengths of XML methods. At the same time, despite the challenging comparison, HGCLR has shown that HTC methods remain relevant and effective in certain scenarios. The results suggest that HTC methods could benefit from adaptation to larger datasets, which would require refinements to cope with increasing complexity.
7 Conclusion
Our experiments show that XML methods are very competitive and should be considered for HTC tasks. In particular, CascadeXML performed well and, on the NYT dataset, outperformed the state-of-the-art HTC methods HGCLR and HBGL. The XML methods have the conceptual advantage that they do not require a pre-defined label hierarchy but instead construct an artificial hierarchy to partition the label space, which makes them applicable to a wider range of multi-label classification problems, including hierarchical text classification as shown here. In the other direction, we found that HTC models are hardly applicable to the XML datasets due to limited scalability. Our findings suggest that future work on HTC should compare to XML methods and that ranking-based metrics P@k and R-Precision should be considered for a more thorough comparison of methods.
Limitations
We do not evaluate the training time of the various models, as the original papers lack sufficient information, and hardware configurations were adjusted during experimentation to suit specific model-dataset combinations. This was particularly necessary for running the HTC models on the XML datasets, in the case of HGCLR, with the help of the authors, but unsuccessful. Nevertheless, we have noted that XML models train much faster than HTC models.
{ack}